# Glued lattices are better quantizers than $K_{12}$

### 01 Dec 2023 v1

@@ I Introduction
In this paper, we prove that $K_{12}$, contrary to the popular belief, is not the optimal lattice quantizer in 12 dimensions. We do this by designing two better lattices and computing their second moments exactly. These new lattices are constructed using gluing theory, which was introduced by Conway and Pless for self-dual binary codes [19] and by Conway and Sloane for integral lattices [20]. Using this theory, self-dual integral lattices were constructed as the union of a finite number of translates of a given base lattice, which is typically a product lattice [21], [1, Sec. 3 in Ch. 4]. Gluing or similar techniques have not, as far as we know, been applied in the quest for good lattice quantizers, which are not necessarily integer lattices. That is the scope of the present paper.
@@ III Gluing Theory
is obtained, where $\{{\boldsymbol{0}}\}\subseteq\Gamma\subseteq L^{*}/L$. We call this construction “gluing.” The term was introduced in [19, 20, 21] for the special case when $L$ is a product lattice or binary product code; we use it more generally for any lattice constructed by (6). In one extreme, (6) yields $L$ and in the other extreme $L^{*}$. Intermediate choices of $\Gamma$ can yield interesting families of lattices or nonlattice packings, such as the Coxeter lattices $A_{n}^{r}$ [23], [13, Sec. 5.1] and $D_{n}^{+}$ [1, p. 119]. The so-called Construction A can be seen as a special case of (6) with $L=\sqrt{2}{\mathbb{Z}}^{n}$ and $\Gamma$ being a rescaled binary block code [1, pp. 137–141, 182–185].
@@ IV-B Glued $D_{6}\times D_{6}$
${\boldsymbol{g}}_{0}=[&0&0&0&0&0&0&],\\ {\boldsymbol{g}}_{1}=\big{[}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\big{]},\\ {\boldsymbol{g}}_{2}=[&1&0&0&0&0&0&],\\ {\boldsymbol{g}}_{3}=\big{[}&-\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\tfrac{1}{2}&\big{]}.\end{array}$ (43)
@@ V Conclusions
As a byproduct, improved lattice quantizers can also be obtained in dimensions $13$ to $15$, since the presently best known lattices in these dimensions are product lattices involving $K_{12}$ [16, Tab. I]. The new best known NSMs in these dimensions are $0.070974$, $0.071399$, and $0.071656$ for $n=13$, $14$, and $15$, respectively.

### 08 Dec 2023 v2

@@ I Introduction
In this paper, we prove that $K_{12}$, contrary to the popular belief, is not the optimal lattice quantizer in 12 dimensions. We do this by designing two better lattices and computing their second moments exactly. These new lattices are constructed using gluing theory, which was introduced by Conway et al. for self-dual block codes in [19] and for integral lattices in [20]. Using this theory, self-dual integral lattices were constructed as the union of a finite number of translates of a given base lattice, which is typically a product lattice [21], [1, Sec. 3 in Ch. 4]. Gluing or similar techniques have not, as far as we know, been applied in the quest for good lattice quantizers, which are not necessarily integer lattices. That is the scope of the present paper.
@@ III Gluing Theory
is obtained, where $\{{\boldsymbol{0}}\}\subseteq\Gamma\subseteq L^{*}/L$. We call this construction “gluing.” The term was introduced in [19, 20, 21] for the special case when $L$ is a product code or product lattice; as in [22], we use it more generally for any lattice constructed by (6). In one extreme, (6) yields $L$ and in the other extreme $L^{*}$. Intermediate choices of $\Gamma$ can yield interesting families of lattices or nonlattice packings, such as the Coxeter lattices $A_{n}^{r}$ [23], [13, Sec. 5.1] and $D_{n}^{+}$ [1, p. 119]. The so-called Construction A can be seen as a special case of (6) with $L=\sqrt{2}{\mathbb{Z}}^{n}$ and $\Gamma$ being a rescaled binary block code [1, pp. 137–141, 182–185].
@@ IV-B Glued $D_{6}\times D_{6}$
${\boldsymbol{g}}_{0}=[&&0&&0&&0&&0&&0&&0&],\\ {\boldsymbol{g}}_{1}=\big{[}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&\big{]},\\ {\boldsymbol{g}}_{2}=[&&1&&0&&0&&0&&0&&0&],\\ {\boldsymbol{g}}_{3}=\big{[}&-&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&&\tfrac{1}{2}&\big{]}.\end{array}$ (43)
@@ V Conclusions
As a byproduct, an improved $13$-dimensional lattice quantizer is obtained. The previously best reported lattice for $n=13$ is a product of $K_{12}$ and a scaled integer lattice $a{\mathbb{Z}}$, with an NSM of $0.071035$ [15, Tab. I]. Replacing $K_{12}$ with the new best $12$-dimensional lattice in a similar product construction yields a slightly improved NSM of $0.070974$.